find two consecutive intergers such that the sum of the first and 3 times the second is 55
Answer by pwac(253)

Let first integer be = x
so second integer is is x+1
We now have to develop an equation that matches the question
So x+3times(x+1)=55 which looks like this:
x+3(x+1)=55 remove bracket by multiplying by the 3
x+3x+3=55 which is 4x+3=55 subtract 3 from both sides
4x=52 divide both sides by 4
x=13
x+1=14
Pete